BlankSpace:

Solve for x: |2x + 6| − 4 = 20 (1 point) x = 9 and x = 11 x = −9 and x = 15 x = 9 and x = −15 No solution

563blackghost:

This includes an absolute so there are two different values. \(\large\bf{(2x+6)-4=20}\) \(\large\bf{-(2x+6)-4=20}\) Simplify the two to find your solutions.

BlankSpace:

it wont load.. please no latex

Marziman:

The answer is C. X = 9 and X = -5

Marziman:

I mean, X = 9 and X = -15

563blackghost:

This includes an absolute so there are two different values. (2x+6)−4=20 −(2x+6)−4=20 Simplify the two to find your solutions.

BlankSpace:

oh okay

563blackghost:

sorry for the latex ;-;

BlankSpace:

latex aint loading. LOL

BlankSpace:

i need help with one more

BlankSpace:

probably one more

563blackghost:

sure